# Geometric Scaling of Bayesian Inference in LLMs

### v2

@@ 6.3. Architectural Trade-offs: Efficiency vs Interpretability
The 4:1 query-to-KV ratio forces keys and values to serve multiple query groups simultaneously, preventing the sharp specialization seen in standard MHA. However, the model compensates by distributing information across dimensions (2D manifolds, distributed attention patterns) while maintaining overall Bayesian computational structure.
The Mistral family provides a complementary negative case: sliding-window attention and MoE routing both weaken dynamic focusing even when static geometry (orthogonality, value manifolds) remains intact.
Implication for deployment: GQA is suitable for production deployment where efficiency matters, but researchers studying mechanisms or building interpretability tools should prefer standard MHA for clearer geometric signatures.
@@ 6.4. Depth Drives Richer Representations
Deeper architectures develop richer uncertainty representations requiring additional dimensions. Crucially, entropy parameterization persists - the additional dimension is not noise but structured geometry enabling more nuanced posterior modeling.
This suggests depth enables transformers to represent multimodal or high-dimensional uncertainty distributions while shallower models compress to 1D entropy coordinates.
@@ 6.5. Layer-wise Functional Specialization
Consistent patterns emerge across models:
Layer 0: Setup phase - high key similarity, moderate attention entropy. Likely initializes representations (position embeddings, token embeddings) before geometric structure emerges.
Layers 1 - N-4: Core computation - strong orthogonality, progressive attention focusing. These layers perform Bayesian hypothesis discrimination and evidence accumulation.
Final 3 - 4 layers: Collapse phase - orthogonality weakens, attention sharpens dramatically. Hypothesis frames collapse as model commits to output distribution.
The Mistral models follow the same structural pattern for value manifolds and key orthogonality, but do not show the expected sharpening of attention in the elimination and refinement stages. This is consistent with constraints on global routing (sliding window) and fragmented updates (MoE).
This functional stratification mirrors the three-stage inference process: binding context $\rightarrow$’ eliminating hypotheses $\rightarrow$’ refining output.
Bayes calibration in natural language vs. wind—-tunnel tasks.
Entropy calibration errors in the ICL setting (0.31–0.44 bits across models) are considerably larger than those observed in the controlled wind—-tunnel experiments (typically below 0.1 bits). This gap is expected: natural—-language prompts introduce substantial semantic ambiguity, and production models are not trained on the synthetic labeling distribution used in our ICL task. The key result is therefore not the absolute calibration error but the systematic correspondence between value—-manifold coordinates, model predictive entropy, and the analytical Bayesian posterior as evidence accumulates. This indicates that the same geometric substrate identified in wind—-tunnel training is actively used by production models during inference.
@@ 6.6. Robustness and Limitations
Domain restriction functions as a natural intervention: it reduces the multiplicity of task-specific inference modes activated by mixed prompts. When prompts come from a single coherent domain (e.g., mathematics), the model operates in a more homogeneous inference regime, revealing the same collapsed 1D manifold observed in wind-tunnel settings. Importantly, domain restriction does not prove that the model performs “true– Bayesian inference on natural language; rather, it shows that a Bayesian geometric coordinate system is embedded in its representation space and can be isolated when the prompt distribution reduces task heterogeneity.
Pythia-410M departs from this pattern, exhibiting an intrinsically low-dimensional value space ($_{1}{+}PC$$_{2}\approx 99.7\%$) under both mixed and mathematics-only prompts (Table Table 1), suggesting that its final-layer values are effectively collapsed regardless of domain.
@@
The emergence of 2D or multi-lobed manifolds in deeper or larger models is not yet theoretically understood. These structures may encode multimodal uncertainty, semantic clustering, training-set heterogeneity, or task-mixture effects. Likewise, the interaction between positional embeddings, local attention kernels, and geometric formation remains an open problem, especially in sliding-window or hybrid transformer—-SSM architectures.
@@
Overall, while the geometric signatures are strikingly robust, a complete causal and mechanistic account of their formation - and of their architectural modulation - remains an important direction for future work.
@@ 7. Discussion
This paper extends the geometric account of Bayesian inference developed in Papers 1—-2 to production-scale language models. Three results stand out: domain restriction collapses value manifolds to the one-dimensional geometry characteristic of exact Bayesian wind-tunnel tasks; transformers navigate this geometry during inference in the SULA experiment; and static geometric signatures - entropy-ordered value manifolds and orthogonal key frames - appear across all architectures evaluated, including GQA, sliding-window, and MoE variants.
@@
Our analysis complements circuit-level studies such as induction heads, copy heads, and pattern-matching mechanisms (Elhage et al., 2021a; Olsson et al., 2022a). Those works identify specialized components; our findings describe the global geometric scaffold in which such components operate. Attention focusing determines which tokens are consulted; key orthogonality creates separable hypothesis directions; and value manifolds encode uncertainty along low-dimensional axes. Mapping specific heads onto regions or branches of this geometry is a natural next step.
@@
The consistency of value manifolds and orthogonal key frames across all models indicates that the representational substrate for Bayesian inference is an architectural invariant. Dynamic focusing, by contrast, depends on routing capacity: full-sequence attention exhibits strong progressive sharpening; GQA reduces it; sliding-window attention and MoE routing weaken or fragment it. This behavior matches the frame—-precision dissociation predicted in Paper 2: the hypothesis frame (keys) stabilizes early and robustly, whereas the precision of posterior refinement is sensitive to architectural constraints.
@@
These findings suggest a geometric foundation for understanding transformer behavior. Value manifolds offer a representation of uncertainty; orthogonal key frames support hypothesis discrimination; and attention focusing provides the mechanism for posterior refinement when architecture permits it. Together, these components form a scalable, architecture-agnostic computational template for approximate Bayesian inference in modern LLMs.
@@ 7.1. Limitations and Future Directions
Our study focuses on dense MHA, GQA, and the first widely deployed combination of sliding-window attention and MoE routing (Mistral). Other architectures - notably state-space models such as Mamba and hybrid transformer—-SSM designs - may require specialized extraction methods, and their geometric structure remains an open question.
@@
We performed forward-pass value interventions on Pythia-410M in the SULA setting to test whether the entropy-aligned value axis is merely correlational or plays a causal role. For each of two “Bayesian” layers (L=12L{=}12 and L=23L{=}23), we first precomputed a unit direction $u_{\text{ent}}$ whose coefficient $v\cdot u_{\text{ent}}$ correlates with predictive entropy. We then applied three families of interventions at that layer, each with a matched random-axis control: (i) axis-cut, which removes the component of $v$ along $u_{\text{ent}}$; (ii) axis-only, which projects $v$ onto the one-dimensional subspace spanned by $u_{\text{ent}}$; and (iii) axis-shift, which adds $\pm 1\sigma$ along $u_{\text{ent}}$ based on the empirical standard deviation of $v\cdot u_{\text{ent}}$.
Representationally, the entropy-aligned axis is clearly special. Cutting along $u_{\text{ent}}$ drives the correlation between $v\cdot u_{\text{ent}}$ and model entropy from $\rho\approx 0.27$$0.32$ down to nearly zero (and sometimes slightly negative), whereas cutting along a random axis of equal dimensionality leaves the correlation largely intact (changes on the order of $10^{-2}$). Conversely, axis-only projections along $u_{\text{ent}}$ preserve or slightly sharpen the correlation, while axis-only projections onto a random axis almost completely destroy it. Axis-shift interventions along $u_{\text{ent}}$ produce small but directional changes in SULA calibration (e.g., $1\sigma$ shifts slightly improving correlation with the Bayesian entropy curve), whereas matched random-axis shifts act as near no-ops on both geometry and calibration.
Behaviorally, however, these single-layer interventions do not yield a clean causal separation. SULA mean absolute error and correlation with the Bayesian posterior change only modestly under both true-axis and random-axis cuts, and the true-axis interventions do not consistently hurt performance more than their random controls. The most conservative interpretation is that the entropy-ordered manifold is a representationally privileged coordinate system for uncertainty, but not a singular bottleneck for Bayesian updating: uncertainty information is likely distributed across multiple dimensions and layers, or the manifold serves as a readout of a more distributed computation. Establishing stronger causal claims will require multi-layer or multi-axis ablations, activation patching, or training-time interventions, which we leave for future work.
@@ 7.2. Related Work
Our analysis connects to several threads in the interpretability and probabilistic—-mechanistic modeling literature. We highlight the relationships most relevant to geometric Bayesian structure.
@@
The constrained-belief-update model of early-layer attention (Aggarwal et al., 2025b) predicts that attention patterns should stabilize early, while finer-grained posterior refinements should occur in value representations. This matches precisely the frame—-precision dissociation predicted in Paper 2 and observed empirically here: keys define a stable hypothesis frame, while values encode uncertainty refinement along low-dimensional axes.
@@
Circuit-level analyses (Elhage et al., 2021b; Olsson et al., 2022b) identify specific mechanisms such as induction heads, pattern-matchers, and copy circuits. Our work operates at a complementary scale: we provide a geometric account of the global representational substrate in which such circuits operate. Attention mechanisms determine which evidence is consulted; key frames define separable hypothesis directions; and value manifolds encode uncertainty along low-dimensional coordinates. Mapping specific circuits onto these global geometric structures is a promising direction for future work.
Overall, the Bayesian geometric lens developed in this series complements prior interpretability approaches by identifying a stable, architecture-spanning substrate for uncertainty representation and by revealing how posterior refinement depends on architectural routing. This perspective helps unify diverse observations in interpretability, probabilistic inference, and model analysis within a single geometric framework.
@@ 8. Conclusion
Transformers trained in tightly controlled settings can implement exact Bayesian inference, and their gradient dynamics generate a low-dimensional geometric substrate that expresses posterior structure. In this work, we asked whether this geometric mechanism survives contact with real language, scale, and heterogeneous training corpora. Our analysis across four model families shows that it does: large language models organize value vectors along a dominant axis that tracks predictive entropy, keys remain close to orthogonal frames, and domain restriction reliably collapses value manifolds into the same low-rank forms observed in synthetic wind-tunnel experiments.
These findings establish that Bayesian-like evidence integration in production models is not a coincidence of sampling or prompt design. It is supported by a persistent geometric invariant—an emergent coordinate system along which uncertainty is expressed and evolves as in-context evidence accumulates. This invariant appears across architectures and training regimes, revealing a structural inductive bias toward representing inference geometrically, even in the absence of any explicit Bayesian objective.
Finally, our causal probes refine the mechanistic picture. Interventions that remove or perturb the entropy-aligned axis selectively disrupt the local geometry of uncertainty, while matched random interventions do not. Yet, these manipulations do not proportionally degrade Bayesian-like behavior, implying that no single direction is solely responsible for the computation. The geometric manifold functions as a stable readout of a distributed inference process rather than a brittle circuit. Understanding how this distributed mechanism arises, and whether it can be shaped, compressed, or accelerated, represents an important frontier for both theory and engineering.
LLMs compute Bayesian updates through distributed mechanisms and inscribe the result onto a low-dimensional entropy-aligned manifold - a representational readout rather than a causal bottleneck.

### v3

@@ 6.3. Architectural Trade-offs: Efficiency vs Interpretability
The 4:1 query-to-KV ratio forces keys and values to serve multiple query groups simultaneously, preventing the sharp specialization seen in standard MHA. However, the model compensates by distributing information across dimensions (2D manifolds, distributed attention patterns) while maintaining overall Bayesian computational structure. The Mistral family provides a complementary case: sliding-window attention weakens dynamic focusing even when static geometry (orthogonality, value manifolds) remains intact. Implication for deployment: GQA is suitable for production deployment where efficiency matters, but researchers studying mechanisms or building interpretability tools should prefer standard MHA for clearer geometric signatures.
@@ 6.4. Depth Drives Richer Representations
Deeper architectures develop richer uncertainty representations requiring additional dimensions. The key point: entropy parameterization persists—the additional dimension is not noise but structured geometry enabling finer posterior modeling. Depth appears to enable transformers to represent multimodal or high-dimensional uncertainty distributions while shallower models compress to 1D entropy coordinates.
@@ 6.5. Layer-wise Functional Specialization
Consistent patterns emerge across models: Layer 0: Setup phase - high key similarity, moderate attention entropy. Likely initializes representations (position embeddings, token embeddings) before geometric structure emerges. Layers 1 - N-4: Core computation - strong orthogonality, progressive attention focusing. These layers perform Bayesian hypothesis discrimination and evidence accumulation. Final 3 - 4 layers: Collapse phase - orthogonality weakens, attention sharpens dramatically. Hypothesis frames collapse as model commits to output distribution. The Mistral models follow the same structural pattern for value manifolds and key orthogonality, but do not show the expected sharpening of attention in the elimination and refinement stages. This is consistent with constraints on global routing (sliding window) and fragmented updates (MoE). This functional stratification mirrors the three-stage inference process: binding context $\rightarrow$ eliminating hypotheses $\rightarrow$ refining output.
Bayes calibration in natural language vs. wind–tunnel tasks.
Entropy calibration errors in the ICL setting (0.31–0.44 bits across models) are considerably larger than those observed in the controlled wind–tunnel experiments (typically below 0.05 bits). This gap is expected: natural–language prompts introduce substantial semantic ambiguity, and production models are not trained on the synthetic labeling distribution used in our ICL task. The key result is therefore not the absolute calibration error but the systematic correspondence between value–manifold coordinates, model predictive entropy, and the analytical Bayesian posterior as evidence accumulates. This indicates that the same geometric substrate identified in wind–tunnel training is actively used by production models during inference.
@@ 6.6. Robustness and Limitations
Domain restriction functions as a natural intervention: it reduces the multiplicity of task-specific inference modes activated by mixed prompts. When prompts come from a single coherent domain (e.g., mathematics), the model operates in a more homogeneous inference regime, revealing the same collapsed 1D manifold observed in wind-tunnel settings. Note that domain restriction does not prove that the model performs “true– Bayesian inference on natural language; rather, it shows that a Bayesian geometric coordinate system is embedded in its representation space and can be isolated when the prompt distribution reduces task heterogeneity. Pythia-410M departs from this pattern, exhibiting an intrinsically low-dimensional value space ($_{1}{+}PC$$_{2}\approx 99.7\%$) under both mixed and mathematics-only prompts (Table 1), suggesting that its final-layer values are effectively collapsed regardless of domain. We attribute this to the Pile’s relatively homogeneous token distribution compared to more diverse web-scale corpora; the model appears to operate in a near-single-mode regime.
@@
The emergence of 2D or multi-lobed manifolds in deeper or larger models is not yet theoretically understood. These structures may encode multimodal uncertainty, semantic clustering, training-set heterogeneity, or task-mixture effects. Likewise, the interaction between positional embeddings, local attention kernels, and geometric formation remains an open problem, especially in sliding-window or hybrid transformer–SSM architectures.
@@ 7. Discussion
This paper extends the geometric account of Bayesian inference developed in Papers I–II to production-scale language models. Three results stand out: domain restriction collapses value manifolds to the one-dimensional geometry characteristic of exact Bayesian wind-tunnel tasks; transformers navigate this geometry during inference in the SULA experiment; and static geometric signatures - entropy-ordered value manifolds and orthogonal key frames - appear across all architectures evaluated, including GQA, sliding-window, and MoE variants.
@@
Our analysis complements circuit-level studies such as induction heads, copy heads, and pattern-matching mechanisms (Elhage et al., 2021a; Olsson et al., 2022b). Those works identify specialized components; our findings describe the global geometric scaffold in which such components operate. Attention focusing determines which tokens are consulted; key orthogonality creates separable hypothesis directions; and value manifolds encode uncertainty along low-dimensional axes. Mapping specific heads onto regions or branches of this geometry is a natural next step.
@@
The consistency of value manifolds and orthogonal key frames across all models indicates that the representational substrate for Bayesian inference is an architectural invariant. Dynamic focusing, by contrast, depends on routing capacity: full-sequence attention exhibits strong progressive sharpening; GQA reduces it; sliding-window attention weakens it by limiting global context access. This behavior matches the frame–precision dissociation predicted in Paper II: the hypothesis frame (keys) stabilizes early and robustly, whereas the precision of posterior refinement is sensitive to architectural constraints.
@@
These findings suggest a geometric foundation for understanding transformer behavior. Value manifolds offer a representation of uncertainty; orthogonal key frames support hypothesis discrimination; and attention focusing provides the mechanism for posterior refinement when architecture permits it. Together, these components form a scalable computational template for approximate Bayesian inference in modern LLMs. Combined with Paper I’s finding that selective state-space models (Mamba) achieve similar Bayesian inference through different routing mechanics, this suggests that content-based value routing—not attention specifically—is the essential architectural ingredient for probabilistic reasoning.
@@ 7.1. Limitations and Future Directions
Our study focuses on dense MHA, GQA, and sliding-window attention (Mistral-7B). Paper I shows that selective state-space models (Mamba) achieve comparable or better Bayesian inference in wind tunnels via content-based routing through input-dependent state selection rather than attention. Extending the geometric extraction methods developed here to Mamba and hybrid transformer–SSM designs—and verifying whether the same entropy-ordered manifolds emerge—is an important direction for unifying the geometric theory across architectures with content-based routing.
@@
We performed forward-pass value interventions on Pythia-410M in the SULA setting to test whether the entropy-aligned value axis is merely correlational or plays a causal role. For each of two “Bayesian” layers (L=12L{=}12 and L=23L{=}23), we first precomputed a unit direction $u_{\text{ent}}$ whose coefficient $v\cdot u_{\text{ent}}$ correlates with predictive entropy. We then applied three families of interventions at that layer, each with a matched random-axis control: (i) axis-cut, which removes the component of $v$ along $u_{\text{ent}}$; (ii) axis-only, which projects $v$ onto the one-dimensional subspace spanned by $u_{\text{ent}}$; and (iii) axis-shift, which adds $\pm 1\sigma$ along $u_{\text{ent}}$ based on the empirical standard deviation of $v\cdot u_{\text{ent}}$. Representationally, the entropy-aligned axis is clearly special. Cutting along $u_{\text{ent}}$ drives the correlation between $v\cdot u_{\text{ent}}$ and model entropy from $\rho\approx 0.27$$0.32$ down to nearly zero (and sometimes slightly negative), whereas cutting along a random axis of equal dimensionality leaves the correlation largely intact (changes on the order of $10^{-2}$). Conversely, axis-only projections along $u_{\text{ent}}$ preserve or slightly sharpen the correlation, while axis-only projections onto a random axis almost completely destroy it. Axis-shift interventions along $u_{\text{ent}}$ produce small but directional changes in SULA calibration (e.g., $1\sigma$ shifts slightly improving correlation with the Bayesian entropy curve), whereas matched random-axis shifts act as near no-ops on both geometry and calibration. Behaviorally, however, these single-layer interventions do not yield a clean causal separation. SULA mean absolute error and correlation with the Bayesian posterior change only modestly under both true-axis and random-axis cuts, and the true-axis interventions do not consistently hurt performance more than their random controls. The most conservative interpretation is that the entropy-ordered manifold is a representationally privileged coordinate system for uncertainty, but not a singular bottleneck for Bayesian updating: uncertainty information is likely distributed across multiple dimensions and layers, or the manifold serves as a readout of a more distributed computation. Establishing stronger causal claims will require multi-layer or multi-axis ablations, activation patching, or training-time interventions, which we leave for future work.
@@
  • Test whether domain restriction generalizes beyond templated prompts to arbitrary single-domain inputs (complex reasoning, mixed-register technical prose) to strengthen the bridge between controlled and naturalistic settings.
@@ 7.2. Related Work
Our analysis connects to several threads in the interpretability and probabilistic–mechanistic modeling literature. We highlight the relationships most relevant to geometric Bayesian structure.
@@
The constrained-belief-update model of early-layer attention (Agarwal et al., 2025) predicts that attention patterns should stabilize early, while finer-grained posterior refinements should occur in value representations. This matches precisely the frame–precision dissociation predicted in Paper II and observed empirically here: keys define a stable hypothesis frame, while values encode uncertainty refinement along low-dimensional axes.
@@
Circuit-level analyses (Elhage et al., 2021b; Olsson et al., 2022a) identify specific mechanisms such as induction heads, pattern-matchers, and copy circuits. Our work operates at a complementary scale: we provide a geometric account of the global representational substrate in which such circuits operate. Attention mechanisms determine which evidence is consulted; key frames define separable hypothesis directions; and value manifolds encode uncertainty along low-dimensional coordinates. Mapping specific circuits onto these global geometric structures is a promising direction for future work. Overall, the Bayesian geometric lens developed in this series complements prior interpretability approaches by identifying a stable, architecture-spanning substrate for uncertainty representation and by revealing how posterior refinement depends on architectural routing. This perspective helps unify diverse observations in interpretability, probabilistic inference, and model analysis within a single geometric framework.
@@ 8. Conclusion
This trilogy establishes a unified account of Bayesian inference in neural sequence models:
  • Paper I shows which architectures can implement Bayesian inference, taxonomized by three inference primitives: belief accumulation, belief transport, and random-access binding. Transformers realize all three; Mamba realizes two; LSTMs realize only accumulation of static sufficient statistics; MLPs realize none.
  • Paper II shows how gradient descent learns to implement these primitives through dynamics exhibiting a mechanistic analogy to EM, sculpting a characteristic geometry: low-dimensional value manifolds, orthogonal key frames, and progressive attention focusing.
  • This paper shows that the geometric substrate for the inference primitives persists at scale. Across four model families, we find the same value-manifold structure, key orthogonality, and domain-specific collapse that wind-tunnel models exhibit.
Our analysis shows that large language models organize value vectors along a dominant axis that tracks predictive entropy, keys remain close to orthogonal frames, and domain restriction reliably collapses value manifolds into the same low-rank forms observed in synthetic experiments. This persistent geometric invariant appears across architectures and training regimes, revealing a structural inductive bias toward representing inference geometrically, even in the absence of any explicit Bayesian objective. Our causal probes refine the mechanistic picture. Interventions that remove or perturb the entropy-aligned axis selectively disrupt the local geometry of uncertainty, while matched random interventions do not. Yet these manipulations do not proportionally degrade Bayesian-like behavior, implying that no single direction is solely responsible for the computation. The geometric manifold functions as a stable readout of a distributed inference process rather than a brittle circuit. The primitives framework from Paper I explains what capabilities each architecture has; the dynamics from Paper II explain how these capabilities are learned; and this paper establishes that the resulting geometry persists even when the training data lacks ground-truth posteriors. Together with controlled wind-tunnel verification of exact Bayesian inference and gradient-level explanations of its emergence, these results complete a three-part account of existence, mechanism, and persistence at scale. Content-based value routing is not merely sufficient for Bayesian inference—it is a stable representational bias induced by modern transformer architectures and training.